Write a question for which the following equation
Answer #1

F=k*q1*q2/r^2

F= F12+F13 =4*10^9*10*10^-9*1*10^-9/r^2 +4*10^9*10*10^-91*10^-9/r^2

Find the force on charge q1(=10^10^-9C) exerted by two equal charges q2 & q3 (=1*10^-9C) separated by distance 'r'

arrangement: q2-------r--------q1----------r---------q3 (question is: combined force on q1 by q1&q2)